Prove that f(x) = is uniformly continuous on (1,00) and not uniformly continuous on (0,1). (19 pts)
Answer #1

22 For x,y 7,1 We have I fex) - fly = 1 $2] = 14-3/4+3)=14-411&+1. or, I fex - fly) = y-x: 2 12-yl 1 ( 1 + (***,871) ay < 212Now, I fex) – flys] = lyde- $2l = 1 m-(m+) = 17€ ands 12-y] = 1 hari Ismai-rml m.mti Immti-rim m.imti weft It my X Test to lt